OpenStudy (anonymous):

what is equal to 2 square root -27 -6i square root 3 6i square root three 6 square root 3i -6 square root 3

OpenStudy (mertsj):

\[2\sqrt{-27}=2\sqrt{9(3)(-1)}=2(3)i \sqrt{3}=6i \sqrt{3}\]